​(d) Find the critical​ t-value that corresponds to

95​% confidence. Assume 13 degrees of freedom.

​(Round to three decimal places as​ needed.)

I do not understand the steps to get the right tail please show all work
Answer #1

Solution :

Given that ,

degrees of freedom = 13

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Using t table,

t /2,df = t 0.025 , 13 = 2.160

t critical value = 2.160